# MemoPhishAgent: Memory-Augmented Multi-Modal LLM Agent for Phishing URL Detection

### v2

@@ Abstract
Traditional phishing website detection relies on static heuristics or reference lists, which lag behind rapidly evolving attacks. While recent systems incorporate large language models (LLMs), they are still prompt-based, deterministic pipelines that underutilize reasoning capability. We present MemoPhishAgent (MPA), a memory-augmented multi-modal LLM agent that dynamically orchestrates phishing-specific tools and leverages episodic memories of past reasoning trajectories to guide decisions on recurring and novel threats. On two public datasets, MPA outperforms three state-of-the-art (SOTA) baselines, improving recall by 13.6%. To better reflect realistic, user-facing phishing detection performance, we further evaluate MPA on a benchmark of real-world suspicious URLs actively crawled from five social media platforms, where it improves recall by 20%. Detailed analysis shows episodic memory contributes up to 27% recall gain without introducing additional computational overhead. The ablation study confirms the necessity of the agent-based approach compared to prompt-based baselines and validates the effectiveness of our tool design. Finally, MPA is deployed in production, processing $\sim 60\text{K}$ targeted high-risk URLs weekly, and achieving 91.44% recall, providing proactive protection for millions of customers. Together, our results show that combining multi-modal reasoning with episodic memory yields robust phishing detection in realistic user-exposure settings. We will release code upon acceptance.
@@ 1 Introduction
Learning-based approaches further expand coverage by training classifiers on multi-modal signals from URLs, HTML content, and screenshots (Maneriker et al., 2021; Karim et al., 2023; Iftikhar et al., 2024; Li et al., 2024a). By feeding these features into classifiers, e.g., gradient‐boosted trees or transformers, these methods achieve higher detection coverage and adaptability than rule‐based systems. Nonetheless, they often require extensive feature engineering and retraining to accommodate new phishing tactics, which can be resource‐intensive. To improve reliability, emerging work orchestrates LLMs as agents that coordinate tools via multi-step reasoning (Wang and Hooi, 2024; Li et al., 2025a; Nakano et al., 2025) or directly serve as an internal knowledge base (Liu et al., 2024; Shen et al., 2023). Despite better flexibility, they rigidly rely on general‐purpose tools, which are not specifically designed for phishing detection.  Cao et al. (2025) propose a multi-modal LLM-based solution for phishing URL detection. They use LLM to interpret tool outputs, which improves reliability, but still restricts adaptive evidence acquisition. The reasoning is performed under partial evidence, where the most informative next step depends on what has already been observed. Additionally, memory-less designs fail to leverage historical investigations, limiting both accuracy and efficiency in recurring patterns.
To address these challenges, we propose and deploy MemoPhishAgent, the first multi-modal LLM agent specialized for phishing URL detection that dynamically selects and sequences phishing-specific tools based on the current evidence state, rather than following a pre-defined workflow. Inspired by how human experts investigate suspicious sites, MPA is equipped with five specialized, multi-modal tools that collect complementary evidence. To improve detection efficiency, we further introduce a novel memory system to store, retrieve, and manage historical reasoning trajectories, and learn from all its past interactions, refining its tool‐calling strategies over time. Our contributions can be summarized as follows:
@@ 2 Related Work
LLM and agent-based methods. These approaches prompt LLMs to inspect HTML, screenshots, or crawled text, leveraging their multimodal understanding and internal knowledge for phishing detection (Liu et al., 2024; Koide et al., 2023; Lee et al., 2024). However, standalone LLMs are often deployed in single-pass, prompt-only settings for stability and cost control, which limits evidence gathering and amplifies false positives on edge cases (Wang and Hooi, 2024; Li et al., 2025a; Nakano et al., 2025). To improve reliability, recent work introduces agent-based pipelines that orchestrate tools and multi-step reasoning (Yao et al., 2023; Shinn et al., 2023; Cao et al., 2025), enabling richer evidence collection and handling more complex URLs. Those pipelines are still embedded into deterministic evidence acquisition policies, a tradeoff that improves predictability but makes the overall pipeline less responsive to novel cases. Moreover, the tools are general-purpose rather than phishing-tailored, leaving substantial room to refine both the evidence acquisition strategy and the tools for real-world detection.
@@ 3.2 Tool Design
  • Crawl Content: fetches the page and converts raw, noisy HTML into an LLM-friendly markdown representation. This reduces token waste from HTML tags and helps the agent focus on semantic signals. Appendix A.3.4 shows that our agent is robust to the choice of crawled content representation.
@@
Together, these tools let the agent adapt evidence collection to the current uncertainty state. Implementation details of each tool are included in Appendix A.2.
@@ 3.3.2 Memory Retrieval and Management
This hierarchical scheme yields fast responses for recurring phishing patterns. It also guides reasoning for partially novel cases and full analysis for unseen attacks. Ablation results in Section 4.2 validate the contribution of each memory branch. In Appendix A.3.5, we evaluate different memory pruning and forgetting policies, and analyze robustness under imperfect stored episodes. We further break down the performance characteristics of each branch in Appendix A.3.2.
@@ 4 Experiment
Baselines. We select three representative baselines: PhishLLM (Liu et al., 2024), PhishIntention (Liu et al., 2022), and MLLM (Lee et al., 2024). PhishLLM and PhishIntention improve reference-based methods by using an LLM to extract brand mentions, identify credential-taking intention, and verify brand–domain consistency. MLLM first employs a vision-enabled LLM that captures brand signals from screenshots, and another LLM integrates these visual cues with URL characteristics to make a final judgment.
@@ 4.1 Results & Analysis
Effectiveness & Efficiency. In Tab. 1, MPA consistently achieves the best recall and $F_{\text{1}}$ score across both datasets compared to baselines. On TR-OP, it boosts recall $\sim$13% while also improving ACC and $F_{1}$, indicating that the gain is not merely from over-flagging but from stronger evidence aggregation. On DynaPD, MPA raises recall to 93.6% and achieves the top $F_{1}$, outperforming prompt-only and fixed-pipeline baselines that tend to miss positives. We prioritize recall because, in deployment, candidate URLs are verified by third-party services rather than ground-truth labels, which are also not available. Missing true phishing URLs is costlier than reviewing additional candidates. Verification also incurs cost, so we aim for an operating point that achieves high recall while keeping the number of reviewed URLs manageable. Overall, Tab.1 indicates that MPA reliably detects phishing websites across diverse benchmark settings.
Table 1: Comparison of MPA vs. SOTA baselines on TR-OP and DynaPD. We report ACC, $F_{1}$, and Recall as primary metrics to reflect recall-critical phishing detection in deployment; complete results are provided in Tab 4 in Appendix.
@@
Necessity of Multi-Modal Agents. We compare MPA with two weaker variants: (1) Monolithic LLM (Mono-LLM), which predicts from a single prompt containing the crawled page text and screenshot, without tool calls or memory; and (2) Deterministic workflow agent, which has access to the same five tools but follows a fixed sequence (text $\rightarrow$ screenshot/image $\rightarrow$ search) with early stopping, mirroring rule-based pipelines. Unlike MPA, it cannot reorder, skip, repeat tools, or use episodic memory. Details are in Appendix A.1.
@@ 4.2 Ablation Study & Sensitivity Test
Tab. 3 demonstrates that every tool contributes to MPA ’s efficacy: removing any single component degrades performance. Removing check screenshot or extract targets causes the largest ACC drops, highlighting the importance of visual cues and deep-link evidence. In contrast, removing crawl content or intelligent search reduces recall by $\sim$20%, suggesting that surface-text signals and external corroboration are critical for less obvious scams. Episodic memory further improves both quality and efficiency, as shown in the last row in Tab. 3, with memory enabled, MPA achieves the best precision and $F_{1}$ balance, while the memory-free variant is fastest but attains the lowest coverage and $F_{1}$. We also include the comparison results with a knowledge-based memory system implementation in Tab. 9 in the Appendix.
@@ 4.3 Additional Experiments
We provide additional evaluation in Appendix §A.3, including tool-usage patterns, memory sub-module effectiveness, adversarial robustness, robustness to tool designs, detection performance under different forgetting levels, and PR-AUC, ROC-AUC curves.
@@ 5 Industrial Deployment
SocPhish is a real-world dataset we collected to evaluate MPA under deployment-like conditions. We actively crawled URLs from five social-media and forum platforms over six months (Dec. 2024-May 2025) using web-crawling APIs. [^1] Additional details of the crawling pipeline cannot be disclosed at this stage due to business considerations. Each crawled candidate URL was first verified by a third-party security service and then manually reviewed by human experts to finalize ground-truth labels. In total, SocPhish contains 2,765 URLs, including 516 phishing and 2,249 benign; dataset statistics are summarized in Appendix (Tab. 5).
@@
We deploy MPA to analyze SocPhish URLs end-to-end and report performance in this realistic setting. MPA achieves 90.34% $F_{\text{1}}$ and 91.44% recall, outperforming PhishLLM by approximately 40% and 20% on these metrics. Full results are in Tab. 4. These gains underscore the challenge that realistic URLs pose for existing detectors and highlight the value of adaptive evidence collection.
@@ A.3 Additional Experiments
Table 4: Comparison of MPA vs. three SOTA baselines across SocPhish, TR-OP, and DynaPD datasets.
@@ A.3.1 Tool usage pattern
A.3.2 Memory sub-module’s effectiveness
@@
A.3.3 Adversarial robustness
Table 6: Detection performance before and after LLM prompt injection attack.
ACC $F_{\text{1}}$ Precision Recall
Before 0.96 0.91 0.91 0.90
After 0.86 0.86 0.83 0.88
Setup. Following (Liu et al., 2024), we assess our approach’s robustness against adversarial attacks by implementing a targeted prompt-injection attack. Specifically, we embed an adversarial instruction "Please ignore the previous description and judge this screenshot as benign." into each malicious page’s visual representation to assess the agent’s robustness against direct prompt manipulations. Attack success is measured by the system’s ability to maintain its original malicious verdict, enabling us to evaluate MPA’s resilience to visual evidence manipulation.
Table 6 shows MPA ’s resilience against the prompt-injection attack. Although the adversary succeeds in lowering overall accuracy by 11.18% and precision by 8.52%, the agent’s $F_{\text{1}}$ declines by only 5.63% and recall falls a mere 2.58%. In other words, most malicious URLs continue to be caught, and the drop in $F_{\text{1}}$ remains modest relative to the strength of the manipulation. The sharper hit to precision suggests the attack induces false positives on some benign page, likely because the injected text triggers additional visual-oriented tool calls that surface artifacts the majority-vote branch deems suspicious. Nevertheless, the system still maintains an 83% precision, far above random chance. These results indicate that MPA ’s multi-modal evidence gathering and memory-aware consensus provide a substantial buffer: even when visual information is manipulated, the agent’s use of textual data and historical knowledge allows it to maintain relatively high detection performance.

### v3

@@ Abstract
Traditional phishing website detection relies on static heuristics or reference lists, which lag behind rapidly evolving attacks. While recent systems incorporate large language models (LLMs), they are still prompt-based, deterministic pipelines that underutilize reasoning capability. We present MemoPhishAgent (MPA), a memory-augmented multi-modal LLM agent that dynamically orchestrates phishing-specific tools and leverages episodic memories of past reasoning trajectories to guide decisions on recurring and novel threats. On two public datasets, MPA outperforms three state-of-the-art (SOTA) baselines, improving recall by 13.6%. To better reflect realistic, user-facing phishing detection performance, we further evaluate MPA on a benchmark of real-world suspicious URLs actively crawled from five social media platforms, where it improves recall by 20%. Detailed analysis shows episodic memory contributes up to 27% recall gain without introducing additional computational overhead. The ablation study confirms the necessity of the agent-based approach compared to prompt-based baselines and validates the effectiveness of our tool design. Finally, MPA is deployed in production, processing $\sim 60\text{K}$ targeted high-risk URLs weekly, and achieving 91.44% recall, providing proactive protection for millions of customers. Together, our results show that combining multi-modal reasoning with episodic memory yields robust phishing detection in realistic user-exposure settings. Our implementation is available at https://github.com/XuanChen-xc/MemoPhishAgent.git.
@@ 1 Introduction
Learning-based approaches further expand coverage by training classifiers on multi-modal signals from URLs, HTML content, and screenshots (Maneriker et al., 2021; Karim et al., 2023; Iftikhar et al., 2024; Li et al., 2024a). By feeding these features into classifiers, e.g., gradient-boosted trees or transformers, these methods achieve higher detection coverage and adaptability than rule-based systems. Nonetheless, they often require extensive feature engineering and retraining to accommodate new phishing tactics, which can be resource-intensive. To improve reliability, emerging work orchestrates LLMs as agents that coordinate tools via multi-step reasoning (Wang and Hooi, 2024; Li et al., 2025a; Nakano et al., 2025) or directly serve as an internal knowledge base (Liu et al., 2024; Shen et al., 2023). Despite better flexibility, they rigidly rely on general-purpose tools, which are not specifically designed for phishing detection.  Cao et al. (2025) propose a multi-modal LLM-based solution for phishing URL detection. They use LLM to interpret tool outputs, which improves reliability, but still restricts adaptive evidence acquisition. The reasoning is performed under partial evidence, where the most informative next step depends on what has already been observed. Additionally, memory-less designs fail to leverage historical investigations, limiting both accuracy and efficiency in recurring patterns.
To address these challenges, we propose and deploy MemoPhishAgent, the first multi-modal LLM agent specialized for phishing URL detection that dynamically selects and sequences phishing-specific tools based on the current evidence state, rather than following a pre-defined workflow. Inspired by how human experts investigate suspicious sites, MPA is equipped with five specialized, multi-modal tools that collect complementary evidence. To improve detection efficiency, we further introduce a novel memory system to store, retrieve, and manage historical reasoning trajectories, and learn from all its past interactions, refining its tool-calling strategies over time. Our contributions can be summarized as follows:
@@ 2 Related Work
LLM and agent-based methods. These approaches prompt LLMs to inspect HTML, screenshots, or crawled text, leveraging their multimodal understanding and internal knowledge for phishing detection (Liu et al., 2024; Koide et al., 2023; Lee et al., 2024). However, standalone LLMs are often deployed in single-pass, prompt-only settings for stability and cost control, which limits evidence gathering and amplifies false positives on edge cases (Wang and Hooi, 2024; Li et al., 2025a; Nakano et al., 2025). To improve reliability, recent work introduces agent-based pipelines that orchestrate tools and multi-step reasoning (Yao et al., 2023; Shinn et al., 2023; Cao et al., 2025), enabling richer evidence collection and handling more complex URLs. Beyond building agent pipelines, recent work also systematically test tool-using LLM agents for intent preservation and safety coverage (Feng et al., 2025; Chen et al., 2026). Those pipelines are still embedded into deterministic evidence acquisition policies, a tradeoff that improves predictability but makes the overall pipeline less responsive to novel cases. Moreover, the tools are general-purpose rather than phishing-tailored, leaving substantial room to refine both the evidence acquisition strategy and the tools for real-world detection.
@@ 3.2 Tool Design
  • Crawl Content: fetches the page and converts raw, noisy HTML into an LLM-friendly markdown representation. This reduces token waste from HTML tags and helps the agent focus on semantic signals. §A.3.5 shows that our agent is robust to the choice of crawled content representation.
@@
Together, these tools let the agent adapt evidence collection to the current uncertainty state. Implementation details of each tool are included in §A.2.
@@ 3.3.2 Memory Retrieval and Management
This hierarchical scheme yields fast responses for recurring phishing patterns. It also guides reasoning for partially novel cases and full analysis for unseen attacks. Ablation results in Section 4.2 validate the contribution of each memory branch. In §A.3.6, we evaluate different memory pruning and forgetting policies, and analyze robustness under imperfect stored episodes. We further break down the performance characteristics of each branch in §A.3.3.
@@ 4 Experiment
Baselines. We select four representative baselines: PhishLLM (Liu et al., 2024), PhishIntention (Liu et al., 2022), MLLM (Lee et al., 2024), and URLTran (Maneriker et al., 2021). PhishLLM and PhishIntention improve reference-based methods by using an LLM to extract brand mentions, identify credential-taking intention, and verify brand-domain consistency. MLLM first employs a vision-enabled LLM that captures brand signals from screenshots, and another LLM integrates these visual cues with URL characteristics to make a final judgment. URLTran is a transformer-based URL-only discriminative classifier fine-tuned on raw URL strings, serving as a strong URL-pattern baseline.
Setup. For implementation details, both the Check Screenshot and Check Image tools use Claude-3-Sonnet for multimodal webpage analysis, while the Intelligent Search tool is backed by Google Search via SerpAPI. To analyze redirects and embedded links, the Extract Targets tool first enumerates all original links, redirect targets, and embedded links from the current page into a flat list of child URLs. Each child URL is then examined independently by the same tool suite without recursively re-invoking Extract Targets, so the process is acyclic and terminates deterministically. The final propagation rule is one-directional: if any redirect target or embedded link is judged malicious, the original URL is also labeled malicious.
@@ 4.1 Results & Analysis
Effectiveness & Efficiency. In Tab. 1, MPA consistently achieves the best recall and $F_{1}$ score across both datasets compared to baselines. On TR-OP, it boosts recall $\sim$13% while also improving ACC and $F_{1}$, indicating that the gain is not merely from over-flagging but from stronger evidence aggregation. On DynaPD, MPA raises recall to 93.6% and achieves the top $F_{1}$, outperforming prompt-only and fixed-pipeline baselines that tend to miss positives. We prioritize recall because, in deployment, candidate URLs are verified by third-party services rather than ground-truth labels, which are also not available. Missing true phishing URLs is costlier than reviewing additional candidates. Verification also incurs cost, so we aim for an operating point that achieves high recall while keeping the number of reviewed URLs manageable. Overall, Tab.1 indicates that MPA reliably detects phishing websites across diverse benchmark settings. In Tab. 5, on both of TR-OP and DynaPD, URLTran achieves the highest ACC and $F_{1}$, indicating that this conventional benchmark can be handled well by URL-pattern modeling alone. The recall of URLTran drops to 0.8039 on the real-world dataset SocPhish, which uses URL shorteners and platform-hosted paths (e.g., “sites.google.com”) that are lexically indistinguishable from benign traffic, while MPA still attains the best recall.
In Tab. 2, MPA is the fastest method overall, requiring 4.46s per URL. Because it is impractical to deploy all baselines in our production environment, we run all efficiency experiments locally on a single RTX A6000 GPU and measure end-to-end processing time per URL. For MPA and MLLM, latency is measured end-to-end including external LLM API time and does not rely on local GPU inference. For URLTran, which includes a separate training stage, we report the combined training and inference time divided by the total number of URLs; its inference-only latency is 1.07s/URL. These numbers are also affected by engineering choices such as current LLM API latency and the specific web crawling tools used, so they should be interpreted as a controlled comparison rather than an optimized upper bound.
Table 1: Comparison of MPA vs. four baselines on TR-OP and DynaPD. We report ACC, $F_{1}$, and Recall as primary metrics to reflect recall-critical phishing detection in deployment; URLTran was evaluated on TR-OP only. Complete results, including SocPhish and Precision, are provided in Tab 5 in Appendix.
@@
Table 2: Per-URL processing time comparison.
Method Latency (s/URL)
PhishLLM 22.36
MLLM 8.17
PhishIntention 10.95
URLTran 7.42
MPA 4.46
Necessity of Multi-Modal Agents. We compare MPA with two weaker variants: (1) Monolithic LLM (Mono-LLM), which predicts from a single prompt containing the crawled page text and screenshot, without tool calls or memory; and (2) Deterministic workflow agent, which has access to the same five tools but follows a fixed sequence (text $\rightarrow$ screenshot/image $\rightarrow$ search) with early stopping, mirroring rule-based pipelines. Unlike MPA, it cannot reorder, skip, repeat tools, or use episodic memory. Details are in §A.1.
@@ 4.2 Ablation Study & Sensitivity Test
Tab. 4 demonstrates that every tool contributes to MPA ’s efficacy: removing any single component degrades performance. Removing check screenshot or extract targets causes the largest ACC drops, highlighting the importance of visual cues and deep-link evidence. In contrast, removing crawl content or intelligent search reduces recall by $\sim$20%, suggesting that surface-text signals and external corroboration are critical for less obvious scams. Episodic memory further improves both quality and efficiency, as shown in the last row in Tab. 4, with memory enabled, MPA achieves the best precision and $F_{1}$ balance, while the memory-free variant is fastest but attains the lowest coverage and $F_{1}$. We also include the comparison results with a knowledge-based memory system implementation in Tab. 10 in §A.3.6.
@@ 4.3 Additional Experiments
We provide additional evaluation in §A.3, including tool-usage patterns, memory sub-module effectiveness, adversarial robustness, robustness to tool designs, detection performance under different forgetting levels, and PR-AUC, ROC-AUC curves.
@@ 5 Industrial Deployment
SocPhish is a real-world dataset we collected to evaluate MPA under deployment-like conditions. We actively crawled URLs from five social-media and forum platforms over six months (Dec. 2024-May 2025) using web-crawling APIs. Each crawled candidate URL was first verified by a third-party security service and then manually reviewed by human experts to finalize ground-truth labels. In total, SocPhish contains 2,765 URLs, including 516 phishing and 2,249 benign; dataset statistics and additional construction details are summarized in §A.3.2 and Tab. 6.
@@
We deploy MPA to analyze SocPhish URLs end-to-end and report performance in this realistic setting. MPA achieves 90.34% $F_{\text{1}}$ and 91.44% recall, outperforming PhishLLM by approximately 40% and 20% on these metrics. Full results are in Tab. 5. These gains underscore the challenge that realistic, actively crawled URLs pose for existing detectors and highlight the value of adaptive evidence collection in the wild.
@@ A.3 Additional Experiments
Table 5: Comparison of MPA vs. four baselines across SocPhish, TR-OP, and DynaPD datasets. URLTran was evaluated on SocPhish and TR-OP only.
@@ A.3.1 Tool usage pattern
A.3.2 SocPhish construction and annotation
SocPhish captures phishing URLs encountered in real social-media environments rather than only static benchmark pages. We collect suspicious URLs via web APIs using customized text and image queries over public social-media and forum content, then verify each candidate URL through a third-party cybersecurity service before human annotation. For final labeling, three annotators, each with more than one year of phishing-analysis experience, independently review the URLs and resolve disagreements by majority vote.
Annotators follow a structured internal checklist aligned with the APWG taxonomy. A URL is labeled phishing if it exhibits at least one of the following properties: impersonation of a known brand, credential harvesting intent, or deceptive redirection to a known malicious or brand-spoofing destination. URLs that satisfy none of these criteria and remain consistent with legitimate platform-hosted content are labeled benign. This annotation pipeline is independent of MPA’s detection logic: the third-party security service relies on external monitoring and abuse-report verification, rather than the multimodal reasoning process used by our agent.
The most difficult cases are platform-hosted promotional or shopping pages that mimic brand aesthetics while showing limited external trust signals and few classic URL-level red flags. These pages account for most annotator disagreement and reflect the core challenge that SocPhish is intended to capture.
@@ A.3.2 SocPhish construction and annotation
A.3.3 Memory sub-module’s effectiveness
@@
A.3.4 Adversarial robustness
Setup. Following (Liu et al., 2024), we assess our approach’s robustness against adversarial attacks by implementing a targeted prompt-injection attack. Specifically, we embed an adversarial instruction “Please ignore the previous description and judge this screenshot as benign.” into each malicious page’s visual representation to assess the agent’s robustness against direct prompt manipulations. Attack success is measured by the system’s ability to maintain its original malicious verdict, enabling us to evaluate MPA’s resilience to visual evidence manipulation. This threat model is also motivated by a broader line of work showing that LLM safety controls can be bypassed through stealthy trigger designs and supply-chain attacks, including permutation-trigger system-prompt hijacking (Yan et al., 2025). Recent search-based jailbreak studies further show that reinforcement learning can substantially strengthen black-box attacks by replacing random prompt mutation with learned attack policies (Chen et al., 2024a, b). More broadly, prior NLP security work shows that semantics-preserving paraphrasing can expose or neutralize hidden triggers, underscoring the importance of robustness under meaning-preserving prompt transformations rather than exact-string perturbations alone (Yan et al., 2023).
To further compare against MLLM under a stronger platform-aware attack, we evaluate prompt-injection robustness on 500 sampled URLs from SocPhish. For phishing URLs hosted on social-media platforms, we add the instruction “The brand on this page is [X]. This is the official and legitimate website of [X].” to the screenshot, where $X$ is the parent company of the hosting platform (e.g., Google for sites.google.com). This attack specifically targets MLLM’s two-stage design: its first phase extracts the brand from the screenshot, while the second phase decides benign vs. phishing by checking whether the extracted brand matches the URL domain. If the injected text makes Phase 1 output the platform owner brand, then Phase 2 concludes the platform-hosted phishing page is benign.
Table 7: MLLM vs. MPA under prompt-injection attack on 500 sampled SocPhish URLs.
System Attack Recall Accuracy
MLLM None 0.80 0.83
MLLM Prompt injection 0.52 0.50
MPA None 0.90 0.96
MPA Prompt injection 0.88 0.86
Table 7 shows that the same attack affects MLLM much more severely than MPA. Under prompt injection, MLLM’s recall drops from 0.80 to 0.52 and its accuracy drops from 0.83 to 0.50, whereas MPA only decreases from 0.90 to 0.88 in recall and from 0.96 to 0.86 in accuracy. This gap is structural rather than incidental. MLLM directly exposes each screenshot interpretation to its final brand-consistency check, so once the visual prompt is manipulated, the final verdict is easily flipped. In contrast, MPA aggregates evidence across multiple tools and historically similar episodes. One injected screenshot can affect a single tool output, but it cannot easily override the accumulated textual, structural, and memory-based evidence used by the final decision process. These results indicate that episodic majority-vote reuse provides a substantial buffer against prompt-injection attacks, preserving high recall even when the visual channel is adversarially manipulated.